OpenStudy (anonymous):

For rules of exponents when dividing...you subtract the exponent in the numerator from the one in the denominator? i.e. x^3/x^4 you would do 4-3? or 3-4?

OpenStudy (shubhamsrg):

x^3 = x*x*x x^4 = x*x*x*x when you divide, what'll be your answer?

OpenStudy (anonymous):

@shubhamsrg thats where im confused...idk if it would be x^-1 or x^1

OpenStudy (shubhamsrg):

(x*x*x)/(x*x*x*x) 3 x's get cancelled you'll be left with 1/x you must be knowing of 1/x = x^-1 or x^1 from this, compare x^(3-4) and x^(4-3) to check which one fits

OpenStudy (anonymous):

you must be knowing of 1/x = x^-1 or x^1 ^^^ how did you get x^-1 or x^1 from 1/x?

OpenStudy (shubhamsrg):

1/x = x^-1 both are the same thing -_-

OpenStudy (shubhamsrg):

anyways, to answer your question directly x^a /x^b = x^(a-b)

OpenStudy (anonymous):

yeah but then you put the x^1 too.

OpenStudy (anonymous):

thanks for all the help!

OpenStudy (shubhamsrg):

what ?

OpenStudy (shubhamsrg):

hmm, welcome